OpenStudy (anonymous):

Write the equation of a line in slope-intercept form passing through (3, 5) and (2, 9)

OpenStudy (anonymous):

as x varies from 2 to 3, y changesfrom 9 to 5 thus the slope is -4. so y = -4x +(y-intercept).

OpenStudy (anonymous):

x=2, y=9, and so each unit along the x axis, y loses 4 units so at x=0 y=9+2*4=17

OpenStudy (anonymous):

y=17-4x